OpenStudy (anonymous):

Suppose that you make the golf team you need to score no more than 74 on average over 5 games. If you scored 77,71,77, and 67 in your first 4 games. What is the highest score you can shoot in your 5th game and still make the team?

OpenStudy (anonymous):

find a value that is divisible by five that equals 74...

OpenStudy (anonymous):

74 = (77+71+77+67+x)/5 solve for x.

OpenStudy (anonymous):

74(5)= 77+71+77+67+x 370=292+x 78=x so 78 is the highest score the team can get to equal 74.

OpenStudy (anonymous):

@Nikebauer12 wouldn't it be more educational to involve the OP with a discussion rather than just solving the problem for them? No one is really interested in whether or not WE can solve it...

OpenStudy (anonymous):

I appreciate everyones help,and i didnt take the answer and run. Isaw how he did it

OpenStudy (anonymous):

@pg.calc well if you show them step by step how to do they will also understand it better.

OpenStudy (anonymous):

hopefully you learned from it. good luck.